# Optimal Approximations for the Requirement Cut Problem on Sparse Graph Classes

### v2

@@ 4.2 Algorithm
In this section we will see how our algorithm can give us a $\mathcal{O}(\log(\sigma(G,\mathcal{S})))$-approximation for the Requirement Cut problem where $\sigma(G,\mathcal{S})$ is the number of distinct minimal Steiner trees in the graph with respect to the given terminal sets. Computing $\sigma(G,\mathcal{S})$ is #$\texttt{\#}P$-Complete is since it is a more general case of counting paths [36]. But, thankfully, we can upper-bound it by $\tau(G)\cdot g$ (1) which is easy to compute through Kirchhoff’s Theorem [30]. When clear from the context, we shall denote the number $\sigma(G,\mathcal{S})$ simply by $\sigma$.
@@
4:Choose $\alpha=\frac{1}{c\cdot\log(\sigma)}$ with $c$ large enough (see Section 4.2)
@@
The rest of this section of the paper will be dedicated to proving this theorem:
@@
In other words the theorem provides $\mathcal{O}(\log(n))$-factor approximation for all graphs described in (Section 3) and others. It also matches the $\mathcal{O}(\log(g))$-approximation on trees given by [25]. We believe that our analysis is also simpler, since it only involves one of the two phases described in [25] (i.e., the second one).
@@
We will call a terminal set $i$ that lies in at least $r_{i}$ components satisfied.
@@
Let $\alpha=\frac{1}{c\cdot\log(\sigma)}$ with $c$ a constant chosen large enough (Section 4.2) and define $x_{e}\sim\text{Uniform}(0,\alpha)$ picked independently for each $e\in E$. Let $d:E\to[0,1]$ be the solution returned by Relaxed LP, and let $\mathcal{T}_{i}$ be the set of all Steiner trees of group $g_{i}$ in $G$. By LP feasibility, we have:
@@
We rely on Equation 1 and the probability that any edge $e$ is cut is given by
@@
This completes the proof of 3.
Now that we have established that each Steiner tree is expected to be cut at least $r_{i}-1$ times, we analyze the probability that every $S_{i}$-Steiner tree is cut $r_{i}-1$ times. Proving that our algorithm returns a solution using Lemma 4.1.
For this analysis, we shall first do a few adjustments into our instance:
@@
The goal of the next steps of the analysis will be to show that the edges in $C_{2}$ cover the residual requirements $r_{i}^{\prime}$ left after the removal of the edges in $C_{1}$ with high probability.
Since we are now only looking at the new graph $G=(V,E\setminus C_{1})$, let us update some of our bounds from before.
@@
By linearity of expectations we get :
$\forall i,\forall t\in\mathcal{T}_{i},\mathds{E}[Y_{t}]=\sum_{e\in t}\mathds{E}[Z_{e}]\geq\sum_{e\in t}\frac{d(e)}{\alpha}\geq\frac{r^{\prime}_{i}-1}{\alpha}$
@@
Now let us bound the probability of one Steiner tree not being satisfied using a Chernoff bound. For convenience, we drop the indexes on the variables since the calculations hold for every Steiner tree $t$ in $G=(V,E\setminus C_{1})$.
We use the following Chernoff bound (the details for some of the steps are available in Section A.1):
Substituting $\delta=1-\frac{r^{\prime}-2}{\mathds{E}[Y_{t}]}$, we derive:
After developping the exponent, we get:
$\exp({-\frac{(\mathds{E}[Y_{t}]-r^{\prime}-2)^{2}}{2\mathds{E}[Y_{t}]}})$
Using the lower bound on $\mathds{E}[Y_{t}]$ to replace it and simplifying:
Finally, we know that $r^{\prime}\geq 2$.
$\leq\exp\left(-\frac{16\alpha^{2}-8\alpha+1}{2\alpha}\right)$
Substitute $\alpha$ for $\frac{1}{c\cdot\log(\sigma)}$:
$=\exp\left(4-\frac{8}{c\cdot\log(\sigma)}-\frac{c\cdot\log(\sigma)}{2}\right)\leq\exp\left(4-\frac{c\cdot\log(\sigma)}{2}\right)$
$=\frac{\exp(4)}{\sigma^{\frac{c}{2}}}$
Since $\sigma$ is $poly(n)$ this shows that for each $t\in\mathcal{T}_{i}$ the probability of not being satisfied is inverse polynomial in $n$.
Under the condition that $\sigma=poly(n)$, this means that $C{}$ is a solution with high probability i.e. $1-\frac{1}{poly(n)}$.
Choosing alpha
Let us now discuss the details of fixing $\alpha$ or rather $c$. For some part of our analysis we need $c\geq 4$ assuming $\log(n)\geq 1$. Technically speaking, one could have a $\log(\sigma)$-approximation even for instances where $\sigma$ is not $poly(n)$. But this would not lead to a “single-log” approximation, as the factor is only $\mathcal{O}(\log(n))$ when $\sigma=poly(n)$ or smaller.
5 Depth of Series Parallel Graphs
In this section, we refine our approximation bounds for the Requirement Cut problem on series-parallel graphs by analyzing differently the algorithm given by Emek and Peleg[15]. This different analysis coupled with the $\mathcal{O}(\log(g))$-approximation for trees [25] yields a $\mathcal{O}(m\cdot\log(g))$-approximation for series-parallel graphs of depth $m$. Series-parallel graphs in general cannot be embedded into trees with less than $\log(n)$ expected distorsion [26]. This means that general series-parallel graphs currently have the same approximation ratio for our problem as general graphs, i.e., $\mathcal{O}(\log(g)\cdot\log(n))$. In fact, the infinite class of series-parallel graphs that Gupta et al. [26] introduce as having distortion no less than $\log(n)$ when embedded into trees have a 0ptof $\mathcal{O}(\log(n))$. Meaning that our result provides a credible alternative also in some cases when the 0ptis smaller than $\log(n)$ and not only when it is a given constant.
Refer to captiondepth $=3$depth $=5$depth $=7$
Figure 5: Series-parallel graphs of 0pt3, 5, and 7. These depths correspond to the number of nested compositions in the graph’s construction and form a subclass of graphs with bounded tree-depth, enabling single-log approximations for Requirement Cut. Note that the number of paths in those graphs can be exponential and that each black line represents an arbitrarily long path in this example.
Additionally, series-parallel graphs of bounded depth are a generalization of another well known class of structures in physics known as melonic graphs [1, 3]. Those graphs are used in modeling condensed matter and black holes through the Sachdev–Ye–Kitaev Model (SYK) and its generalization the Carrozza, Tanasa, Klebanov, and Tarnopolsky (CTKT) model. Melonic graphs are built recursively using an operation called “bananification” ([1] did not invent it but uses it), this operation creates series-parallel graphs with some 0ptthat corresponds to the order of the melonic graph.
On the other hand, the depth parameter that we consider captures exactly the depth of the tree decomposition of the given series-parallel graph. This draws a parallel to the well-studied structural parameter tree-depth [33], which can be equivalently seen to capture the width of the tree decomposition together with its depth. As series-parallel graphs are of treewidth 2, series-parallel graphs of bounded depth are thus special cases of graphs of bounded tree-depth.
In 2006, Emek and Peleg [15] presented an algorithm that probabilistically embeds series-parallel graphs into distributions over their spanning trees with expected distortion $\mathcal{O}(\log(n))$. Their approach constructs a spanning tree recursively from the composition trace of the graph, ensuring that, in expectation, distances between any pair of vertices remain within a logarithmic factor of the original graph.
We modify the analysis by focusing on the 0ptof the series-parallel graph, in other words, it is the depth of its composition tree as defined in [15]. Formally, this is the number of times the compositions in the trace alternate between Series and Parallel. Note that the 0ptof a series-parallel graph can grow linearly in the worst case (it is upper bounded by $\frac{n}{2}$. However, our claim is not to have a better approximation for all series-parallel graphs but simply that for any fixed 0pt, all series-parallel graph of that 0pthave an efficient approximation. As we discussed in the introduction, even series-parallel graphs of depth $\mathcal{O}(\log n)$ pose a serious obstacle towards better approximation for Requirement Cut.
[…]
Before giving the proof and our statement, it is good to give some context and explanations regarding Emek and Peleg’s algorithm. Their algorithm is called Construct_Tree. Construct_Tree takes as input a series-parallel graph $G$ with two designated terminal vertices $x$ and $y$. The algorithm returns a spanning tree $T$ of $G$. The goal is to ensure that for any pair of vertices, the expected stretch is small between the shortest distance in $G$ and the distance in $T$ is small.
The algorithm operates recursively, using the composition tree (or decomposition trace) of $G$, which reflects how the graph was built from smaller series-parallel components via series and parallel composition. It handles these two cases differently:
  • Series composition: If $G$ is obtained by gluing smaller graphs $G_{1},\ldots,G_{k}$ in series, the algorithm recursively constructs a spanning tree $T_{j}$ for each $G_{j}$, then glues them together in series to form $T$. This process does not increase stretch for any pair $(u,v)$ contained in any $G_{i}$.
  • Parallel composition: If $G$ offers multiple subgraphs $G_{1},\ldots,G_{k}$ that each connect $x$ to $y$, the algorithm first builds a spanning tree $T_{j}$ for each $G_{j}$. It identifies the shortest $(x,y)$-path $P_{k^{*}}$ among them, and keeps $T_{k^{*}}$ untouched. For the other $T_{j}$ (with $j\neq k$), it removes one edge chosen uniformly at random from the path $P_{j}$ connecting $x$ and $y$ in $T_{j}$. This breaks the cycles and ensures the union of all modified trees still forms a spanning tree of $G$. This step is the one that might increase the stretch for some $(u,v)$ contained in a $G_{i}$.
The number of compositions is bounded by $m$ since it is the 0ptof the composition trace that we use to compose our graph. To measure the stretch, Emek and Peleg introduce several variables: $\{a_{i}\}_{i=1}^{m}$, $\{b_{i}\}_{i=1}^{m}$ and $d$ positive integers satisfying:
  • $b_{1}=1$,
  • $b_{i+1}\leq a_{i}+b_{i}$ for all $1\leq i<m$,
  • $c_{i}=\sum_{j=1}^{i-1}a_{i}$,
  • $d\leq a_{m}+b_{m}$,
  • $\sum_{i=1}^{m}a_{i}\leq n$.
Intuitively those variables each mean something different in the construction with respect to a pair of vertices $u,v\in G_{j}$, a graph that does not contain $P_{k}$ the shortest $(x,y)$-path. During a parallel composition operation, $d(u,v)$ might increase if an edge in the shortest $(u,v)$-path is removed:
  • $b_{i}$ is the length of the overlap between the shortest $(u,v)$-path and the $(x,y)$-path in $G_{j}$, $P_{j}$. This is the size of the "risk" portion of $P_{j}$.
  • $a_{i}$ is the length of the rest of $P_{j}$.
  • $c_{i}$ is $d(u,v)-b_{i}$, it is what will remain of the $d(u,v)$ no matter what.
  • $d$ is simply $d(x,y)$.
We will now show how to bound the expected distorsion of Emek and Peleg’s algorithm by an $\mathcal{O}(m)$ factor instead of a $\mathcal{O}(\log(n))$.
Theorem 5.1.
Let $\mathcal{T}$ be the output of Algorithm Construct_Tree [15] when invoked on an series-parallel graph G. Then $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(m),\forall(u,v)\in E(G)$. Where $m$ is the depth of the composition trace of $G$.
Then, we have for $\mathds{E}[dist_{\mathcal{T}}(u,v)]$, the distorsion ([15] Lemma 4.1):
In their paper [15], Emek and Peleg show that $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(\log n)$ (Lemma 4.2), meaning that the expected stretch of each edge is at most logarithmic in $n$. That is:
We adopt the same notations as them and with a different analysis obtain a different result depending on $m$: $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(m),\forall(u,v)\in E(G)$
We first rewrite it using $b_{i+1}\leq bi+ai$ which implies that $b_{i}\leq 1+\sum_{j=1}^{i-1}a_{i}$
We replace $c_{i}$, using $c_{i}=a_{1}+...+a_{i-1}$
$=\frac{a_{1}}{a_{1}+1}+\sum\limits_{i=2}^{m}\frac{a_{i}}{1+\sum_{j=1}^{i}a_{j}}\left(\frac{1+2\cdot\sum_{j=1}^{i-1}a_{j}}{1+\sum_{j=1}^{i-1}a_{j}}\right)+\left(\frac{d_{m}+\sum_{i=1}^{m}a_{i}}{1+\sum_{i=1}^{m}a_{i}}\right)$
And now we upper bound $d_{m}$ by using $d_{m}\leq 1+\sum_{i=1}^{m}a_{i}$
$\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq 2m+2$
This series of calculations show that for a given 0pt, all series-parallel graphs of that depth can be embedded into a distribution of spanning trees that give a constant expected distorsion for any pair of vertices $(u,v)$. This fact is interesting on its own theoretically and in our case, we can use it to get an $\mathcal{O}(\log(g))$-approximation for the Requirement Cut problem on series-parallel graphs of bounded 0pt.
6 Conclusion
The Requirement Cut problem generalizes numerous classical cut problems, and has resisted single-logarithmic approximations on general graphs due to unavoidable lower bounds like the cost of tree embeddings. In this work, we identify two structural parameters—the number of (minimal) Steiner trees and the depth of series-parallel graphs—that allow us to break the “double-log” barrier.
Our main takeaway is that graphs with a polynomial number of spanning trees, or more generally instances with a polynomial number of minimal Steiner trees, admit efficient LP-rounding algorithms with approximation ratio $\mathcal{O}(\log(n))$. Additionally series-parallel graphs of bounded depth can be approximated in $\mathcal{O}(\log(g))$, extending known results for trees to a broader class. These findings broaden the landscape of graphs for which single-log approximations are possible which was previously limited to trees.
We believe our work represents a natural limit of improvement (up to constant factors) for $LP$ rounding techniques based on randomized thresholding. Future progress may require fundamentally different approaches. Nonetheless, our results suggest that graphs with polynomially many spanning trees and bounded-depth constructions behave in a “tree-like” manner, opening the door to improved approximation algorithms for other connectivity problems in such graphs.
Appendix A Appendix
A.1 Chernoff Bound Details
We take the Chernoff bound presented in Lemma 4.6 and detail each step here:
@@ A.1 Chernoff Bound Details
Using this bound to upper bound the original expression:
Note: This upper-bounding is possible because the function $f(x)=\exp\left(-\frac{(x-r^{\prime}-2)^{2}}{2x}\right)$ is decreasing for any $x\geq r^{\prime}-2$ which is the case for our case.
@@
$=\exp\left(4-\frac{8}{c\cdot\log(\sigma)}-\frac{c\cdot\log(\sigma)}{2}\right)\leq\exp\left(4-\frac{c\cdot\log(\sigma)}{2}\right)$
$=\frac{\exp(4)}{\sigma^{\frac{c}{2}}}$
This achieves the calculations. You can join back the proof and discussion on $\alpha$ (Section 4.2).

### v3

@@ 4.2 Algorithm
In this section we will see how our algorithm can give us a $\mathcal{O}(\log(\sigma(G,\mathcal{S})))$-approximation for the Requirement Cut problem where $\sigma(G,\mathcal{S})$ is the number of distinct minimal Steiner trees in the graph with respect to the given terminal sets. Computing $\sigma(G,\mathcal{S})$ is $\texttt{\#}P$-Complete since it is a more general case of counting paths which is $\texttt{\#}P$-Complete [valiant_complexity_1979]. However, we can upper-bound it by $\tau(G)\cdot g$ (Claim 1) which is easy to compute through Kirchhoff’s Theorem [kirchhoff_ueber_1847]. When clear from the context, we shall denote the number $\sigma(G,\mathcal{S})$ simply by $\sigma$.
@@
4:Choose $\alpha=\frac{1}{c\cdot\log(\sigma)}$ with $c$ large enough (see Choosing alpha)
@@
The rest of this section of the paper will be dedicated to proving the following theorem:
@@
The algorithm provides an $\mathcal{O}(\log(n))$-factor approximation for all graphs with $\tau(G)=\mathcal{O}(n^{c})$ (see Section 3 for examples). It also matches the $\mathcal{O}(\log(g))$-approximation on trees given by [gupta_improved_2010], while having a simpler process, as it only involves one of their two phases.
@@
Let $\alpha=\frac{1}{c\cdot\log(\sigma)}$ with $c$ a constant chosen large enough (Choosing alpha) and define $x_{e}\sim\text{Uniform}(0,\alpha)$ picked independently for each $e\in E$. Let $d:E\to[0,1]$ be the solution returned by Relaxed LP, and let $\mathcal{T}_{i}$ be the set of all Steiner trees of group $g_{i}$ in $G$. We call a terminal set $i$ that lies in at least $r_{i}$ components satisfied. By LP feasibility, we have:
@@
We rely on eq. 1 and the probability that any edge $e$ is cut is given by
@@
This completes the proof of this Claim.
Now that we have established that each Steiner tree is expected to be cut at least $r_{i}-1$ times, we analyze the probability that every $S_{i}$-Steiner tree is cut $r_{i}-1$ times. Proving that our algorithm returns a solution using Lemma 4.1.For this analysis, we shall first do a few adjustments into our instance:
@@
The goal of the next steps of the analysis will be to show that the edges in $C_{2}$ cover the residual requirements $r_{i}^{\prime}$ left after the removal of the edges in $C_{1}$ with high probability.Since we are now only looking at the new graph $G=(V,E\setminus C_{1})$, let us update some of our bounds from before.
@@
By linearity of expectations we get : $\forall i,\forall t\in\mathcal{T}_{i},\mathds{E}[Y_{t}]=\sum_{e\in t}\mathds{E}[Z_{e}]\geq\sum_{e\in t}\frac{d(e)}{\alpha}\geq\frac{r^{\prime}_{i}-1}{\alpha}$. Now, let us bound the probability of one Steiner tree not being satisfied using a Chernoff bound. We shall drop the indices on the variables since the calculations hold for every Steiner tree $t$ in $G=(V,E\setminus C_{1})$.We use the following Chernoff bound:
@@
Using this bound to upper bound the original expression:Note: This upper-bounding is possible because the function $f(x)=\exp\left(-\frac{(x-r^{\prime}-2)^{2}}{2x}\right)$ is decreasing for any $x\geq r^{\prime}-2$ which is the case for our case.
@@
$\exp\left(4-\frac{8}{c\cdot\log(\sigma)}-\frac{c\cdot\log(\sigma)}{2}\right)\leq\exp\left(4-\frac{c\cdot\log(\sigma)}{2}\right)$
$\frac{\exp(4)}{\sigma^{\frac{c}{2}}}$
Under the condition that $\sigma=poly(n)$, this means that CC{} is a solution with high probability i.e. $1-\frac{1}{poly(n)}$.This concludes the proof of Lemma 4.6
This also completes the proof of Theorem 4.3.
Choosing alpha
Let us now discuss the details of fixing $\alpha=\frac{1}{c\cdot\log(\sigma)}$. For some part of our analysis (see long version of proof of Lemma 4.6) we need $c\geq 4$ assuming $\log(n)\geq 1$.Technically speaking, one could have a $\log(\sigma)$-approximation even for instances where $\sigma$ is not $poly(n)$. But this would not lead to a “single-log” approximation, as the factor is only $\mathcal{O}(\log(n))$ when $\sigma=poly(n)$ or smaller.
5 Depth of Series Parallel Graphs
In this section, we refine our approximation bounds for the Requirement Cut problem on series-parallel graphs by analyzing differently the algorithm given by Emek and Peleg [emek_tight_2010]. This different analysis coupled with the $\mathcal{O}(\log(g))$-approximation for trees [gupta_improved_2010] yields a $\mathcal{O}(m\cdot\log(g))$-approximation for series-parallel graphs of depth $m$. Series-parallel graphs in general cannot be embedded into trees with less than $\log(n)$ expected distortion [gupta_cuts_2004]. This means that general series-parallel graphs currently have the same approximation ratio for our problem as general graphs, i.e., $\mathcal{O}(\log(g)\cdot\log(n))$. In fact, the infinite class of series-parallel graphs that Gupta et al. [gupta_cuts_2004] introduce as having distortion no less than $\log(n)$ when embedded into trees have a 0ptof $\mathcal{O}(\log(n))$. Meaning that our result provides a credible alternative also in some cases when the 0ptis smaller than $\log(n)$ and not only when it is a given constant.
Refer to captiondepth =3=3depth =5=5depth =7=7
Figure 6: Series-parallel graphs of 0pt3, 5, and 7. These depths correspond to the number of nested compositions in the graph’s construction and form an infinite subclass of graphs with bounded tree-depth. Note that the number of paths in those graphs can be exponential. Here, each black line represents an arbitrarily long path.
Additionally, series-parallel graphs of bounded depth are a generalization of another well known class of structures in physics known as melonic graphs [aluffi_motives_2022, baratin_melonic_2014]. Those graphs are used in modeling condensed matter and black holes through the Sachdev–Ye–Kitaev Model (SYK) and its generalization the Carrozza, Tanasa, Klebanov, and Tarnopolsky (CTKT) model. Melonic graphs are built recursively using an operation called “bananification” ([aluffi_motives_2022] did not invent it but uses it), this operation creates series-parallel graphs with some 0ptthat corresponds to the order of the melonic graph.The depth parameter that we consider captures exactly the depth of the tree decomposition of the given series-parallel graph. This draws a parallel to the well-studied structural parameter tree-depth [nesetril_bounded_2012], which can be equivalently seen to capture the width of the tree decomposition together with its depth. As series-parallel graphs are of treewidth 2, series-parallel graphs of bounded depth are thus a special case of graphs of bounded tree-depth.In 2006, Emek and Peleg [emek_tight_2010] presented an algorithm that probabilistically embeds series-parallel graphs into distributions over their spanning trees with expected distortion $\mathcal{O}(\log(n))$. Their approach constructs a spanning tree recursively from the composition trace of the graph, ensuring that, in expectation, distances between any pair of vertices remain within a logarithmic factor of the original graph.We modify the analysis by focusing on the depth of the series-parallel graph. Note that the depth of a series-parallel graph can grow linearly in the worst case (it is upper bounded by $\frac{n}{2}$). However, our claim is not to have a better approximation for all series-parallel graphs but simply that for any fixed depth, all series-parallel graph of that depth have an efficient approximation. As discussed in the introduction, series-parallel graphs of depth $\mathcal{O}(\log n)$ pose a serious obstacle towards better approximation for Requirement Cut since they are as hard as general graphs to approximate.In other words, our analysis proves that fixing the depth at a constant $m$ for series-parallel allows $\mathcal{O}(\log(g))$ for the Requirement Cut problem on this class of graphs.Before giving our theorem and proving it, it is good to give some context regarding Emek and Peleg’s algorithm Construct_Tree. It takes as input a series-parallel graph $G$ with two terminal vertices $x$ and $y$ and returns a spanning tree $T$ of $G$. The goal is to minimize the expected distortion between the shortest distances in $G$ and in $T$.The algorithm operates recursively, using the composition tree (or decomposition trace) of $G$, which reflects how the graph was built from smaller series-parallel components via series and parallel composition. It handles these two cases differently:
  • Series composition: If $G$ is obtained by gluing $G_{1},\ldots,G_{k}$ in series, the algorithm recursively constructs a spanning tree $T_{j}$ for each $G_{j}$, then glues $T_{1},\ldots,T_{k}$ in series to form $T$. This does not increase distortion for any pair $(u,v)$ contained in any $G_{i}$.
  • Parallel composition: If $G$ is obtained by taking $G_{1},\ldots,G_{k}$ that each connect $x$ to $y$, the algorithm first builds a spanning tree $T_{j}$ for each $G_{j}$. It identifies a shortest $(x,y)$-path $P_{k^{*}}$ among them, and keeps $T_{k^{*}}$ untouched. For the other $T_{j}$ (with $j\neq k^{*}$), it removes one edge chosen uniformly at random from the path $P_{j}$ connecting $x$ and $y$ in $T_{j}$. This breaks the cycles and ensures the union of all modified trees still forms a spanning tree of $G$. This step might increase the distortion for some $(u,v)$ contained in any $G_{i}$.
The number of compositions is bounded by $m$ since it is the maximum number of compositions we do. To measure the distortion, Emek and Peleg introduce several variables: $\{a_{i}\}_{i=1}^{m}$, $\{b_{i}\}_{i=1}^{m}$ and $d$ positive integers satisfying: $b_{1}=1$, $b_{i+1}\leq a_{i}+b_{i}$ for all $1\leq i<m$, $c_{i}=\sum_{j=1}^{i-1}a_{i}$, $d\leq a_{m}+b_{m}$, $\sum_{i=1}^{m}a_{i}\leq n$. Intuitively, these variables each mean something different in the construction with respect to a pair of vertices $u,v\in G_{j}$, a graph that does not contain $P_{k}$ the shortest $(x,y)$-path. During a parallel composition operation, $d(u,v)$ might increase if an edge in the shortest $(u,v)$-path is removed:
  • $b_{i}$ is the length of the overlap between the shortest $(u,v)$-path and the $(x,y)$-path in $G_{j}$, $P_{j}$. This is the size of the ”risk“ portion of $P_{j}$.
  • $a_{i}$ is the length of the rest of $P_{j}$.
  • $c_{i}$ is $d(u,v)-b_{i}$, it is what will remain of the $d(u,v)$ no matter what.
  • $d$ is simply $d(x,y)$.
We will now show how to bound the expected distortion of Emek and Peleg’s algorithm by an $\mathcal{O}(m)$ factor instead of a $\mathcal{O}(\log(n))$.
Theorem 5.1.
Let $\mathcal{T}$ be the output of Algorithm Construct_Tree [emek_tight_2010] when invoked on a series-parallel graph G with $m$ the depth of the composition trace of $G$. Then $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(m),\forall(u,v)\in E(G)$.
Proof 5.2.
Let $\mathds{E}[dist_{\mathcal{T}}(u,v)]$, be the distortion between $u,v\in V$ as defined in [emek_tight_2010, Lemma 4.1] :
In their paper [emek_tight_2010], Emek and Peleg show that $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(\log n)$ (Lemma 4.2), meaning that the expected distortion of each edge is at most logarithmic in $n$. That is:
We adopt the same notations as them and with a different analysis obtain a different result depending on $m$: $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(m),\forall(u,v)\in E(G)$We first rewrite it using $b_{i+1}\leq b_{i}+a_{i}$ which implies that $b_{i}\leq 1+\sum_{j=1}^{i-1}a_{i}$
We replace $c_{i}$, using $c_{i}=a_{1}+...+a_{i-1}$
$\frac{a_{1}}{a_{1}+1}+\sum\limits_{i=2}^{m}\frac{a_{i}}{1+\sum_{j=1}^{i}a_{j}}\left(\frac{1+2\cdot\sum_{j=1}^{i-1}a_{j}}{1+\sum_{j=1}^{i-1}a_{j}}\right)+\left(\frac{d_{m}+\sum_{i=1}^{m}a_{i}}{1+\sum_{i=1}^{m}a_{i}}\right)$
And now we upper bound $d_{m}$ by using $d_{m}\leq 1+\sum_{i=1}^{m}a_{i}$
$\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq 2m+2$
This series of calculations show that for a given 0pt, all series-parallel graphs of that depth can be embedded into a distribution of spanning trees that give a constant expected distortion for any pair of vertices $(u,v)$. This fact is interesting on its own theoretically and in our case, we can use it to get an $\mathcal{O}(\log(g))$-approximation for the Requirement Cut problem on series-parallel graphs of bounded 0pt.
6 Conclusion
The Requirement Cut problem generalizes many classical cut problems, and has resisted single-log approximations on general graphs due to unavoidable lower bounds like the cost of tree embeddings. In this work, we identify two structural parameters—the number of (minimal) Steiner trees and the depth of series-parallel graphs—allowing us to break the “double-log” barrier.Our main takeaway is that graphs with a polynomial number of spanning trees, or more generally instances with a polynomial number of minimal Steiner trees, admit efficient LP-rounding algorithms with approximation ratio $\mathcal{O}(\log(n))$. Additionally series-parallel graphs of bounded depth can be approximated in $\mathcal{O}(\log(g))$, extending known results for trees to a broader class.We believe our work represents a natural limit of improvement (up to constant factors) for both our methods. On one hand, $LP$ rounding techniques based on randomized thresholding have been extensively studied and our improvement implicitly defines and calls for graphs of polynomial number of spanning trees which have not been studied before. On the other hand our linear bound in the depth $m$ of series-parallel graphs is likely optimal up to constant factors, since the series-parallel graphs constructed in [gupta_cuts_2004] which exhibit distortion $\Omega(\log n)$ when embedded into trees have depth $\Theta(\log n)$ matching the $\log(n)\cdot\log(g)$-approximation [nagarajan_approximation_2010, gupta_improved_2010] on these graphs. Therefore, any future progress may require fundamentally different approaches.Nonetheless, our results suggest that graphs with polynomially many spanning trees and bounded-depth constructions behave in a “tree-like” manner and allow for single-log approximations for the Requirement Cut problem, opening the door to improved approximation algorithms for other connectivity problems in such graphs.